Find the orthogonal canonical reduction of the quadratic form
1
Expert's answer
2020-02-24T12:25:16-0500

Find the orthogonal canonical reduction of the quadratic form "-x^2+y^2+z^2-6xy-6xz+2yz"

The matrix of the quadratic form


"A=\begin{pmatrix}
 -1 & -3 & -3 \\
 -3 & 1 & 1\\
 -3 & 1 & 1
\end{pmatrix}"

Find the eigenvalues


"A-\lambda I=\begin{pmatrix}
 -1-\lambda & -3 & -3 \\
 -3 & 1-\lambda & 1\\
 -3 & 1 & 1-\lambda
\end{pmatrix}"

The characteristic equation is "det (A-\lambda I)=0"

"A-\lambda I=\begin{vmatrix}
 -1-\lambda & -3 & -3 \\
 -3 & 1-\lambda & 1\\
 -3 & 1 & 1-\lambda
\end{vmatrix}=""=( -1-\lambda)\begin{vmatrix}
 1-\lambda & 1\\
 1 & 1-\lambda
\end{vmatrix}-(-3)\begin{vmatrix}
 -3 & 1\\
 -3 & 1-\lambda
\end{vmatrix}+""+(-3)\begin{vmatrix}
 -3 & 1-\lambda \\
 -3 & 1 
\end{vmatrix}=(-1-\lambda)((1-\lambda)^2-1)+""+3(-3(1-\lambda)+3)-3(-3+3(1-\lambda))=0"

"2\lambda-\lambda^2+2\lambda^2-\lambda^3+9\lambda+9\lambda=0""\lambda(-\lambda^2+\lambda+20)=0""-\lambda(\lambda+4)(\lambda-5)=0"

"\lambda_1=-4"

"\lambda_2=0"

"\lambda_3=5"

Find the eigenvectors

"\lambda=-4"


"\begin{pmatrix}
 -1-\lambda & -3 & -3 \\
 -3 & 1-\lambda & 1\\
 -3 & 1 & 1-\lambda
\end{pmatrix}=\begin{pmatrix}
 3 & -3 & -3 \\
 -3 & 5 & 1\\
 -3 & 1 & 5
\end{pmatrix}"

Perform row operations to obtain the rref of the matrix:


"\begin{pmatrix}
 3 & -3 & -3 \\
 -3 & 5 & 1\\
 -3 & 1 & 5
\end{pmatrix}\sim\begin{pmatrix}
 1 & 0 & -2 \\
 0 & 1 & -1\\
 0 & 0 & 0
\end{pmatrix}"

Solve the matrix equation


"\begin{pmatrix}
 1 & 0 & -2 \\
 0 & 1 & -1\\
 0 & 0 & 0
\end{pmatrix}\begin{pmatrix}
 v_1 \\
 v_2 \\
 v_3 
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\
 0 
\end{pmatrix}"

If we take "v_3=t," then "v_1=2t, v_2=t, v_3=t."

Therefore


"\textbf{v}=\begin{pmatrix}
 2t \\
 t \\
 t 
\end{pmatrix}=\begin{pmatrix}
 2 \\
 1 \\
 1 
\end{pmatrix} t"

The principal axis is

"{1 \over \sqrt{6}}\begin{pmatrix}
 2 \\
 1 \\
 1
\end{pmatrix}"

"\lambda=0"


"\begin{pmatrix}
 -1-\lambda & -3 & -3 \\
 -3 & 1-\lambda & 1\\
 -3 & 1 & 1-\lambda
\end{pmatrix}=\begin{pmatrix}
 -1 & -3 & -3 \\
 -3 & 1 & 1\\
 -3 & 1 & 1
\end{pmatrix}"

Perform row operations to obtain the rref of the matrix:


"\begin{pmatrix}
 -1 & -3 & -3 \\
 -3 & 1 & 1\\
 -3 & 1 & 1
\end{pmatrix}=\begin{pmatrix}
 1 &0 & 0 \\
 0 & 1 & 1\\
 0 & 0 & 0
\end{pmatrix}"

Solve the matrix equation


"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 1\\
 0 & 0 & 0
\end{pmatrix}\begin{pmatrix}
 v_1 \\
 v_2 \\
 v_3 
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\
 0 
\end{pmatrix}"

If we take "v_3=t," then "v_1=0, v_2=-t, v_3=t."

Therefore


"\textbf{v}=\begin{pmatrix}
 0 \\
 -t \\
 t 
\end{pmatrix}=\begin{pmatrix}
 0 \\
 -1 \\
 1 
\end{pmatrix} t"

The principal axis is


"{1 \over \sqrt{2}}\begin{pmatrix}
 0 \\
 -1 \\
 1
\end{pmatrix}"

"\lambda=5"


"\begin{pmatrix}
 -1-\lambda & -3 & -3 \\
 -3 & 1-\lambda & 1\\
 -3 & 1 & 1-\lambda
\end{pmatrix}=\begin{pmatrix}
 -6 & -3 & -3 \\
 -3 & -4 & 1\\
 -3 & 1 & -4
\end{pmatrix}"

Perform row operations to obtain the rref of the matrix:


"\begin{pmatrix}
 -6 & -3 & -3 \\
 -3 & -4 & 1\\
 -3 & 1 & -4
\end{pmatrix}\sim\begin{pmatrix}
 1 & 0 & 1 \\
 0 & 1 & -1\\
 0 & 0 & 0
\end{pmatrix}"

Solve the matrix equation


"\begin{pmatrix}
 1 & 0 & 1 \\
 0 & 1 & -1\\
 0 & 0 & 0
\end{pmatrix}\begin{pmatrix}
 v_1 \\
 v_2 \\
 v_3 
\end{pmatrix}=\begin{pmatrix}
 0 \\
 0 \\
 0 
\end{pmatrix}"

If we take "v_3=t," then "v_1=-t, v_2=t, v_3=t."

Therefore


"\textbf{v}=\begin{pmatrix}
 -t \\
 t \\
 t 
\end{pmatrix}=\begin{pmatrix}
 -1 \\
 1 \\
 1 
\end{pmatrix} t"

The principal axis is


"{1 \over \sqrt{3}}\begin{pmatrix}
 -1 \\
 1 \\
 1
\end{pmatrix}"

"P=\begin{pmatrix}
 2/\sqrt{6} & 0 & -1/\sqrt{3} \\
 1/\sqrt{6} & -1/\sqrt{2} & 1/\sqrt{3}\\
 1/\sqrt{6} & 1/\sqrt{2} & 1/\sqrt{3}
\end{pmatrix}"

"P^{-1}=P^T=\begin{pmatrix}
 2/\sqrt{6} & 1/\sqrt{6} & 1/\sqrt{6} \\
 0 & -1/\sqrt{2} & 1/\sqrt{2}\\
 -1/\sqrt{3} & 1/\sqrt{3} & 1/\sqrt{3}
\end{pmatrix}"

"D=\begin{pmatrix}
 -4 & 0 & 0 \\
 0 & 0 &0\\
 0 & 0 & 5
\end{pmatrix}"

"\begin{pmatrix}
 x_1 \\
 y_1 \\
 z_1
\end{pmatrix}=P^T\begin{pmatrix}
 x \\
 y \\
 z
\end{pmatrix}"

"x_1= (2/\sqrt{6})x+( 1/\sqrt{6})y+( 1/\sqrt{6})z""y_1= (0)x+( -1/\sqrt{2})y+( 1/\sqrt{2})z""z_1= (-1/\sqrt{3})x+( 1/\sqrt{3})y+( 1/\sqrt{3})z"

The orthogonal canonical reduction of the quadratic form is


"-4x_1^2+5z_1^2=0"

The principal axes are


"p_1={1 \over \sqrt{6}}\begin{pmatrix}
 2 \\
 1 \\
 1
\end{pmatrix}, p_2={1 \over \sqrt{2}}\begin{pmatrix}
 0 \\
 -1 \\
 1
\end{pmatrix}, p_3={1 \over \sqrt{3}}\begin{pmatrix}
 -1 \\
 1 \\
 1
\end{pmatrix}"